OpenStudy (anonymous):

Find the slope-intercept form for the line satifying the following conditions. x-intercept 2, y-intercept 1/3

OpenStudy (ajprincess):

Let the slope-intercept form be y=mx+c. When u put x=0 in u y=mx+c u get y-intercept. When u put y=0 in u y=mx+c u get x-intercept. Can u find nw @lawme106?

OpenStudy (anonymous):

1/3=m(0)+c 0=m(2)+c is this the correct form? i'm lost.

OpenStudy (ajprincess):

c=1/3  Getting it?